v=0, u=15m/s1, a=1m/s2
How much displacement and time
1
Expert's answer
2020-06-01T14:25:22-0400

Solution.

"v=0;"

"u=15m\/s;"

"a=1m\/s^2;"

"S-?; t-?;"

"S=\dfrac{u^2-v^2}{2a};"

"S=\dfrac{(15m/s)^2-0}{2\sdot1m/s}=112.5m;"

"a=\dfrac{u-v}{t} \implies t=\dfrac{u-v}{a};"

"t=\dfrac{15m/s-0}{1m/s^2}=15s;"

Answer: "S=112.5m;"

"t=15s."